# Deep Submillimeter and Radio Observations in the SSA22 Field. IV. Spectral Energy Distributions, Star Formation Histories, and the Infrared-Radio Correlation of the 850 -selected SMGs

### v1

@@ Abstract
We analyze the spectral energy distributions (SEDs), star formation histories (SFHs), and infrared-radio correlation (IRRC) of 221 850 -selected submillimeter galaxies (SMGs) in the SSA22 deep field. The average SEDs reveal that SMGs have cooler dust temperatures than local ULIRGs or starburst systems, and that higher-redshift SMGs tend to be more luminous and exhibit warmer dust. The median mass-weighted age is 567 Myr. Most galaxies in our sample began forming 1.68 Gyr after the Big Bang, entered the ‘SMG phase’ after 1 Gyr of evolution—when they are predominantly observed—and largely transitioned out of the ‘SMG phase’ to become quiescent within an additional 0.2 Gyr. A subset of massive galaxies shows rapid early assembly with high star formation efficiencies (0.2-0.8). The majority of SMGs reside at the high-mass end of the star-forming main sequence, with a characteristic stellar mass of M⊙, above which galaxies are predominantly either on the main sequence or already quenched. We observe a downsizing trend: more massive galaxies tend to “mature” earlier, completing their major episodes of star formation at higher redshifts compared to lower-mass systems. Our sample contributes 21% (28%) to the cosmic star formation rate density (stellar mass density), including the overdensity, with its relative contribution peaking at 50-60% in the redshift range . We suggest that 850 surveys may miss a population of faint, warm galaxies at . The median infrared-radio correlation parameter is 2.37, evolving as , likely due to AGN contributions at high redshift and intrinsic differences between low- and high-redshift populations.
@@ I Introduction
In the local universe, ultra-luminous infrared galaxies (ULIRGs) are almost exclusively merger-driven systems (Lonsdale et al., 2006; Geach et al., 2017). By analogy, it was long assumed that high-redshift SMGs are similarly triggered by gas-rich major mergers (Toft et al., 2014), consistent with their often asymmetric morphologies, clumpy structures, and kinematically disordered gas motions (Alaghband-Zadeh et al., 2012; Menéndez-Delmestre et al., 2013). However, alternative models propose that SMGs can be sustained over longer timescales through steady cold gas accretion (Narayanan et al., 2015; McAlpine et al., 2019). Indeed, spatially resolved spectroscopy has revealed rotationally supported disks in many SMGs (Genzel et al., 2011; Swinbank et al., 2011; De Breuck et al., 2014; Barro et al., 2017), and their optical/far-infrared morphologies are often well described by exponential disk profiles (Targett et al., 2013; Hodge et al., 2016; Fujimoto et al., 2018). Recent high-resolution JWST observations further suggest that many SMGs may be massive, isolated disk galaxies whose intense star formation is driven by gravitational instabilities in gas-rich disks (Gillman et al., 2023, 2024; Le Bail et al., 2024; Chan et al., 2025; Hodge et al., 2025; Zhang et al., 2025). Cosmological simulations also support the possibility of SMG-like systems evolving in isolation (Davé et al., 2010, 2011).
@@
An open and intriguing question remains: could a non-negligible fraction of SMGs retain their disk structure and evolve into present-day massive disk galaxies? Observational and theoretical studies suggest that galaxies undergoing major merger events that dramatically alter their structure can subsequently reform disk-like configurations. Simulations indicate that if the merging system remains gas-rich, a rotationally supported disk can re-emerge within a few Gyrs (Barnes, 2002; Springel and Hernquist, 2005; Robertson et al., 2006; Hopkins et al., 2009). Moreover, Governato et al. (2009) emphasized the role of environment, showing that even gas-poor merger remnants can regenerate disks through the accretion of cold streams from the cosmic web or cooling flows from hot halo gas. The model of Barro et al. (2013) also supports this view, suggesting that galaxies can rebuild disk structures via gas-poor mergers or dynamic instabilities prior to final quenching. Additionally, Gillman et al. (2024) found that the stellar disks of SMGs can remain intact and persist until at least 1. An equally intriguing question is whether SMGs that avoid major disruptive events might retain their disk morphology to the present day. Future studies of massive quiescent discs at 1-2 will be essential to address this possibility and refine our understanding of SMG long-term evolutionary pathways.
@@
The persistence of the IRRC at high redshifts presents a theoretical challenge, given the rapidly increasing energy density of the cosmic microwave background (CMB) with redshift () (Murphy, 2009). Under these conditions, inverse Compton (IC) scattering and ionization losses should dominate cosmic-ray (CR) electron cooling, surpassing synchrotron radiation. This could potentially breakdown the IRRC at high redhsift, or equivalently, increase with redshift (Magnelli et al., 2015). However, observations indicate that IC cooling has not yet overtaken synchrotron cooling at least up to 2 (Magnelli et al., 2015). Our sample demonstrates that the IRRC persists at least out to 4, and Delhaize et al. (2017) report that galaxies continue to follow the IRRC even at higher redshifts ( 6). What mechanisms preserve the IRRC at high redshifts? Theoretical models suggest that enhanced galactic magnetic fields could increase the efficiency of synchrotron cooling, compensating for the increased energy losses from IC scattering and ionization due to the stronger CMB (Sargent et al., 2010b; Yoon, 2024). However, magnetic field amplification alone may be insufficient, as it would require extremely high field strengths. Moreover, under such conditions, CR electrons primarily lose energy via ionization and bremsstrahlung rather than synchrotron emission, which would further suppress non-thermal radio emission. Alternative explanations include the possibility that AGN activity dominates the radio emission at high redshifts, thereby maintaining the observed IRRC (Murphy et al., 2009), and/or enhanced free-free (bremsstrahlung) emission at high redshifts supplement the total radio luminosity (Murphy et al., 2009; Magnelli et al., 2015; Delhaize et al., 2017). A combination of these mechanisms is likely responsible for maintaining the correlation in the early universe.
@@ II Data
We present the deepest 850 imaging of the SSA22 deep field to date, combining data from the SCUBA-2 Cosmology Legacy Survey (S2CLS; Geach et al., 2017) with additional observations of the central region (Ao et al., 2017). The total integration time amounts to 91 hours, covering an area of 0.34 deg2, with a central sensitivity reaching 0.79 mJy beam-1. The JVLA 3 GHz observations of the central area of SSA22 field were conducted in 2015 and 2016, achieving a sensitivity of 1.5 Jy beam-1 (before primary beam correction) and an angular resolution of 23 ×20. For detailed descriptions of the SCUBA-2 850 observations and data reduction, we refer readers to Paper II of this series (Zeng et al., 2024); for the radio observations and data processing, see Paper I (Ao et al., 2017) and III.
@@ III.1 Average SED
Assuming a simple blackbody model, we estimate the characteristic dust temperature () from the peak wavelength of the average SED using Wien’s displacement law (Uematsu et al., 2024). When using flux densities in units of [mJy] measured at discrete frequencies, the peak temperature is given by [K] = 5099/. Arp220 and M82 exhibit emission peaks at 70-90 , corresponding to 60-70 K. In contrast, the SSA22 and AS2UDS samples, along with the sample from Michałowski et al. (2010), exhibit cooler dust temperatures around 50 K ( 100 ). The ALESS sample falls between these values, with a temperature of approximately 60 K.
@@ III.2 SED Variations Across Bins
For the results presented above and in the following sections, we should bear in mind several potential biases that may affect their interpretation. First, SED fitting relies on FIR band constraints. This introduces a selection bias toward intrinsically brighter sources at high redshifts, while low-luminosity galaxies at similar redshifts may fall below the FIR detection limit. Consequently, some studies argue that the apparent luminosity evolution of SMGs is largely an artifact of this observational bias (Swinbank et al., 2014; Dudzevičiūtė et al., 2020). However, other works contend that such biases alone cannot fully account for the observed trends (Simpson et al., 2017; Zavala et al., 2018; Lim et al., 2020b). They suggest that infrared luminosity and dust temperature are still governed by underlying physical processes, with dust temperatures evolving with cosmic time (Symeonidis et al., 2013; Sommovigo et al., 2022; Viero et al., 2022; Mitsuhashi et al., 2024). In particular, a more fundamental correlation likely exists among infrared luminosity, dust temperature, and redshift (Symeonidis et al., 2013; Zavala et al., 2018). Second, bandpass selection effects play a significant role (see also Section IV.1 and Section IV.4). Observations at longer (sub)millimeter wavelengths are more sensitive to high-redshift dusty galaxies (Casey et al., 2014; Zavala et al., 2014; Wang et al., 2017), while 850 surveys preferentially select galaxies with cooler dust temperatures (Chen et al., 2021). As a result, 850 surveys are inherently incomplete for warm-dust systems and for galaxies at both low and very high redshifts. Moreover, uncertainties in the detection and physical characterization of high-redshift galaxies, combined with the intrinsic diversity of SMG populations, further complicate the interpretation. A full resolution of these issues lies well beyond the scope of this study. Readers should simply be aware of these potential biases when evaluating our conclusions.
@@
This paints an evolutionary picture: as intense star formation proceeds, large amounts of dust are produced, UV photons from young stars are absorbed, and dust thermal emission strengthens. Coupled with increasing dust column density, the efficiency of dust radiation improves. Generally, galaxies with more vigorous star formation exhibit stronger infrared emission and hotter dust temperatures. However, at a fixed star formation rate, as dust accumulates, the efficiency of heating the outer dust layers decreases, leading to lower infrared luminosity and gradually cooler average dust temperatures. This process also enhances submillimeter emission.
@@ III.3 The radiation Composition of SMGs
The long-wavelength SED is governed by synchrotron radiation from relativistic electrons accelerated in supernova remnants of massive ( 8 M⊙) stars (Condon, 1992). The power-law energy distribution of these electrons produces a corresponding power-law synchrotron spectrum with typical indices of 0.3-2.0. Studies of star-forming galaxies reveal a median spectral index of 0.8 at rest-frame 1.3-10 GHz, flattening at lower frequencies due to free-free absorption and spectral aging from energy losses (An et al., 2021, 2024). Both the high- and low-frequency of the radio spectrum become slightly steeper with increasing stellar mass, a trend that can be explained by age-related synchrotron losses, where cosmic-ray electrons lose energy over time (An et al., 2021). Massive stars have short lifetimes ( 3 107 yr), and the lifetimes of relativistic electrons are also limited ( yr), making radio emission a tracer of relatively recent star formation activity (Condon, 1992). However, accurately observing and constraining the star formation activity in galaxies remains challenging. Although synchrotron radiation is highly directional due to the Doppler effect of relativistic electrons, the observed emission in galaxies typically originates from relatively “old” electrons ( yr), which have long since propagated far ( kpc) from their short-lived ( yr) supernova remnants. As a result, the original sites of electron acceleration have already disappeared, and their spatial distribution has been significantly smoothed out. Consequently, radio emission observed in galaxies appears diffuse across the entire galactic extent (Condon, 1992).
@@
AGN contribute across multiple wavebands, predominantly in the mid-infrared and radio (with additional contributions in the UV and optical for type-1 AGN). The AGN fraction in SMGs can reach 20-40% or higher (Alexander et al., 2005; Laird et al., 2010; Johnson et al., 2013; An et al., 2019; Stach et al., 2019). However, studies have shown that at least 85% of the far-infrared emission in SMGs is powered by star formation activity (Laird et al., 2010). In general, AGN contribute little to the total luminosity and even less to the far-infrared and millimeter emission of SMGs, typically 10% (Johnson et al., 2013). As seen in our results, these findings are well supported.
@@ IV.1.2 Specific Star Formation Rate
The inverse of sSFR corresponds to the characteristic stellar mass doubling timescale (Madau and Dickinson, 2014), i.e., the time required to assemble the current stellar mass at the present SFR. Juneau et al. (2005) used the average timescale () to characterize the transition from burst to quiescent star formation mode across galaxy populations, defining galaxies with shorter than the Hubble time as actively star-forming, and those with longer timescales as transitioning into quiescence. As shown in right panel of Figure 6, applying this criterion yields conclusions consistent with our earlier classification, despite the different threshold. Comparing with the results of Juneau et al. (2005), at similar masses, even the reddest galaxies selected in the optical and near-infrared bands quench more rapidly than SMGs. This highlights that at intermediate to low redshifts (), the majority of cosmic star formation occurs obscured by dust (Bourne et al., 2017; Dunlop et al., 2017; Zavala et al., 2021; Bouwens et al., 2022), and particularly at , dusty star-forming galaxies contribute approximately 80% of the total cosmic star formation activity (Zavala et al., 2021).

### v2

@@ Abstract
We analyze the spectral energy distributions (SEDs), star formation histories (SFHs), and infrared-radio correlation (IRRC) of 221 850 -selected submillimeter galaxies (SMGs) in the SSA22 deep field. The average SEDs reveal that SMGs have cooler dust temperatures than local ULIRGs or starburst systems, and that higher-redshift SMGs tend to be more luminous and exhibit warmer dust. The median mass-weighted age is 567 Myr. Most galaxies in our sample began forming 1.68 Gyr after the Big Bang, entered the ‘SMG phase’ after 1 Gyr of evolution—when they are predominantly observed—and largely transitioned out of the ‘SMG phase’ to become quiescent within an additional 0.2 Gyr. A subset of massive galaxies shows rapid early assembly with high star formation efficiencies (0.2-0.8). The majority of SMGs reside at the high-mass end of the star-forming main sequence, with a characteristic stellar mass of M⊙, above which galaxies are predominantly either on the main sequence or already quenched. We observe a downsizing trend: more massive galaxies tend to “mature” earlier, completing their major episodes of star formation at higher redshifts compared to lower-mass systems. Our sample contributes 21% (28%) to the cosmic star formation rate density (stellar mass density), including the overdensity, with its relative contribution peaking at 50-60% in the redshift range . The median infrared-radio correlation parameter is 2.37, evolving as , likely due to AGN contributions at high redshift and intrinsic differences between low- and high-redshift populations.
@@ I Introduction
In the local universe, ultra-luminous infrared galaxies (ULIRGs) are almost exclusively merger-driven systems (Sanders et al., 1988; Lonsdale et al., 2006). By analogy, it was long assumed that high-redshift SMGs are similarly triggered by gas-rich major mergers (Tacconi et al., 2008), consistent with their often asymmetric morphologies, clumpy structures, and kinematically disordered gas motions (Alaghband-Zadeh et al., 2012; Menéndez-Delmestre et al., 2013). However, alternative models propose that SMGs can be sustained over longer timescales through steady cold gas accretion (Narayanan et al., 2015; McAlpine et al., 2019). Indeed, spatially resolved spectroscopy has revealed rotationally supported disks in many SMGs (Genzel et al., 2011; Swinbank et al., 2011; De Breuck et al., 2014; Barro et al., 2017; Amvrosiadis et al., 2024; Birkin et al., 2024), and their optical/far-infrared morphologies are often well described by exponential disk profiles (Targett et al., 2013; Hodge et al., 2016; Fujimoto et al., 2018). Recent high-resolution JWST observations further suggest that many SMGs/DSFGs may be massive, isolated disk galaxies whose intense star formation is driven by gravitational instabilities in gas-rich disks (Gillman et al., 2023, 2024; Le Bail et al., 2024; Chan et al., 2025; Hodge et al., 2025; Zhang et al., 2025). Cosmological simulations also support the possibility of SMG-like systems evolving in isolation (Davé et al., 2010, 2011).
@@
In the local Universe, galaxies with stellar masses M⊙ are predominantly early-type systems (Kelvin et al., 2014; Moffett et al., 2016; Driver et al., 2022). If we assume that SMGs have gas fractions close to unity and that all gas is eventually converted into stars, most would exceed this mass threshold by 0. But does this necessarily mean that all SMGs evolve into massive elliptical galaxies in the local Universe? Observational and theoretical evidence suggests a more nuanced picture. From a dynamical perspective, Cappellari et al. (2011) showed that about two-thirds of local early-type galaxies (Hubble types S0 and E) are still rotation-dominated, rather than pressure-supported classical ellipticals. Cappellari (2016) further demonstrated that massive local galaxies with log 10.5–11.3 populate distinct structural sequences but remain largely rotation-supported; only the most massive systems (log 11.3) exhibit the kinematic and morphological properties of traditional ellipticals. Gillman et al. (2023) found that, in rest-frame optical imaging, the non-merging SMGs typically exhibits disk-like morphologies (see their Figure 7). Simulations further indicate that even after undergoing extreme major merger events, galaxies can still gradually develop a rotating disk structure provided the merger system remain gas-rich (Barnes, 2002; Springel and Hernquist, 2005; Robertson et al., 2006; Hopkins et al., 2009). Even in gas-poor mergers, the accretion of cold flows from the cosmic web or the cooling of hot halo gas may still facilitate the regrowth of a disk (Governato et al., 2009). We remain open regarding whether all SMGs will eventually evolve into elliptical galaxies. A comprehensive study of massive star-forming and quenching galaxies at both high and low redshifts will be essential to further clarify this evolutionary picture.
@@
The persistence of the IRRC at high redshifts presents a theoretical challenge, given the rapidly increasing energy density of the cosmic microwave background (CMB) with redshift () (Murphy, 2009). Under these conditions, inverse Compton (IC) scattering and ionization losses should dominate cosmic-ray (CR) electron cooling, surpassing synchrotron radiation. This could potentially breakdown the IRRC at high redshift, or equivalently, increase with redshift (Magnelli et al., 2015). However, observations indicate that IC cooling has not yet overtaken synchrotron cooling at least up to 2 (Magnelli et al., 2015). Our sample demonstrates that the IRRC persists at least out to 4, and Delhaize et al. (2017) report that galaxies continue to follow the IRRC even at higher redshifts ( 6). What mechanisms preserve the IRRC at high redshifts? Theoretical models suggest that enhanced galactic magnetic fields could increase the efficiency of synchrotron cooling, compensating for the increased energy losses from IC scattering and ionization due to the stronger CMB (Sargent et al., 2010b; Yoon, 2024). However, magnetic field amplification alone may be insufficient, as it would require extremely high field strengths. Moreover, under such conditions, CR electrons primarily lose energy via ionization and bremsstrahlung rather than synchrotron emission, which would further suppress non-thermal radio emission. Alternative explanations include the possibility that AGN activity dominates the radio emission at high redshifts, thereby maintaining the observed IRRC (Murphy et al., 2009), and/or enhanced free-free (bremsstrahlung) emission at high redshifts supplement the total radio luminosity (Murphy et al., 2009; Magnelli et al., 2015; Delhaize et al., 2017). A combination of these mechanisms is likely responsible for maintaining the correlation in the early universe.
@@ II Data
We present the deepest 850 imaging of the SSA22 deep field to date, combining data from the SCUBA-2 Cosmology Legacy Survey (S2CLS; Geach et al., 2017) with additional observations of the central region (Ao et al., 2017). The total integration time amounts to 91 hours, covering an area of 0.34 deg2, with a central sensitivity reaching 0.79 mJy beam-1. The JVLA 3 GHz observations of the central area of SSA22 field were conducted in 2015 and 2016, achieving a sensitivity of 1.5 Jy beam-1 (before primary beam correction) and an angular resolution of 23 ×20. For detailed descriptions of the SCUBA-2 850 observations and data reduction, we refer readers to Paper II of this series (Zeng et al., 2024); for the radio observations and data processing, see Paper I (Ao et al., 2017) and Paper III.
@@ III.1 Average SED
We select a conservative subsample consisting of sources with redshift 1 and secure counterparts at VLA 3 GHz or MIPS 24 (see Paper III). The average SED derived from these more reliable identifications shows a sharper far-infrared peak, and remains consistent with the average SED of the full sample. This demonstrates that, even if the full sample contains some level of contamination, the ensemble-averaged SED is sufficiently robust and does not alter our main conclusions.
@@
Assuming a simple blackbody model, we estimate the characteristic dust temperature () from the peak wavelength of the average SED using Wien’s displacement law (Uematsu et al., 2024). When using flux densities in units of [mJy] measured at discrete frequencies, the peak temperature is given by [K] = 5099/. Arp220 and M82 exhibit emission peaks at 70-90 , corresponding to 60-70 K. In contrast, the SSA22 and AS2UDS samples exhibit cooler dust temperatures around 50 K ( 100 ). The ALESS sample falls between these values, with a temperature of approximately 60 K.
@@ III.2 SED Variations Across Bins
For the results presented above and in the following sections, we should bear in mind several potential biases that may affect their interpretation. First, SED fitting relies on FIR band constraints. This introduces a selection bias toward intrinsically brighter sources at high redshifts, while low-luminosity galaxies at similar redshifts may fall below the FIR detection limit. Consequently, some studies argue that the apparent luminosity evolution of SMGs is largely an artifact of this observational bias (Swinbank et al., 2014; Dudzevičiūtė et al., 2020; Drew and Casey, 2022). However, other works contend that such biases alone cannot fully account for the observed trends (Zavala et al., 2018; Lim et al., 2020b). They suggest that infrared luminosity and dust temperature are still governed by underlying physical processes, with dust temperatures evolving with cosmic time (Symeonidis et al., 2013; Sommovigo et al., 2022; Viero et al., 2022; Mitsuhashi et al., 2024). In particular, a more fundamental correlation likely exists among infrared luminosity, dust temperature, and redshift (Symeonidis et al., 2013; Zavala et al., 2018). Second, there is a significant bandpass selection effect (see also Section IV.1 and Section IV.4). The selection effect at 850 is fundamentally tied to dust mass. At a fixed infrared luminosity, SMGs/DSFGs with larger dust masses exhibit colder dust SEDs (Cunha et al., 2015). Consequently, an 850 survey with a fixed flux limit may miss sources that are warmer, more luminous in the infrared, but lower in dust mass. This implies that single-band 850 surveys may suffer from incompleteness. Moreover, due to the negative -correction in the submillimeter/millimeter regime, observations at longer wavelengths are more efficient at detecting high-redshift sources (Casey et al., 2014; Zavala et al., 2014; Wang et al., 2017), resulting in samples with higher median redshifts. Adding further complexity are uncertainties in the detection and physical property estimation of high-redshift galaxies, as well as the intrinsic diversity of SMGs themselves—in terms of dust temperature, geometry, star formation histories, and AGN contributions. These issues are beyond the scope of this study. Readers should simply be aware of these potential biases when interpreting our results.
@@
This paints an evolutionary picture: as intense star formation proceeds, large amounts of dust are produced and the dust column density increases, UV photons from young stars are increasingly absorbed, and dust thermal emission strengthens. Galaxies with more vigorous star formation generally exhibit stronger infrared emission. However, at fixed infrared luminosity, an increase in dust mass leads to a larger reservoir of dust to absorb the same radiation field, resulting in a lower equilibrium dust temperature. This cooling effect, combined with higher dust mass, typically enhances emission at submillimeter wavelengths, where the flux density scales approximately as in the Rayleigh–Jeans regime.
@@ III.3 The Radiation Composition of SMGs
The long-wavelength SED is governed by synchrotron radiation from relativistic electrons accelerated in supernova remnants of massive ( 8 M⊙) stars (Condon, 1992). The power-law energy distribution of these electrons produces a corresponding power-law synchrotron spectrum with typical indices of 0.3-2.0. Studies of star-forming galaxies reveal a median spectral index of 0.8 at rest-frame 1.3-10 GHz, flattening at lower frequencies due to free-free absorption and spectral aging from energy losses (Thomson et al., 2019; An et al., 2021, 2024). Both the high- and low-frequency of the radio spectrum become slightly steeper with increasing stellar mass, a trend that can be explained by age-related synchrotron losses, where cosmic-ray electrons lose energy over time (An et al., 2021). Massive stars have short lifetimes ( 3 107 yr), and the lifetimes of relativistic electrons are also limited ( yr), making radio emission a tracer of relatively recent star formation activity (Condon, 1992). However, accurately observing and constraining the star formation activity in galaxies remains challenging. Although synchrotron radiation is highly directional due to the Doppler effect of relativistic electrons, the observed emission in galaxies typically originates from relatively “old” electrons ( yr), which have long since propagated far ( kpc) from their short-lived ( yr) supernova remnants. As a result, the original sites of electron acceleration have already disappeared, and their spatial distribution has been significantly smoothed out. Consequently, radio emission observed in galaxies appears diffuse across the entire galactic extent (Condon, 1992).
@@
AGN contribute across multiple wavebands, predominantly in the mid-infrared and radio (with additional contributions in the UV and optical for type-1 AGN). AGN activity is frequently observed in SMGs, with reported fractions of 15–20% (Wang et al., 2013; An et al., 2019; Stach et al., 2019; Uematsu et al., 2025), brighter and more massive SMGs appear to be more likely to host AGNs (Wang et al., 2013; Cowie et al., 2018; Lim et al., 2020b). However, studies have shown that at least 85% of the far-infrared emission in SMGs is powered by star formation activity (Laird et al., 2010). In general, AGN contribute little to the total luminosity and even less to the far-infrared and millimeter emission of SMGs, typically 10% (Johnson et al., 2013). As seen in our results, these findings are well supported.
@@ IV.1.2 Specific Star Formation Rate
The inverse of sSFR corresponds to the characteristic stellar mass doubling timescale (Madau and Dickinson, 2014), i.e., the time required to assemble the current stellar mass at the present SFR. Juneau et al. (2005) used the average timescale () to characterize the transition from burst to quiescent star formation mode across galaxy populations, defining galaxies with shorter than the Hubble time as actively star-forming, and those with longer timescales as transitioning into quiescence. As shown in right panel of Figure 6, applying this criterion yields conclusions consistent with our earlier classification, despite the different threshold. Comparing with the results of Juneau et al. (2005), at similar masses, even the reddest galaxies selected in the optical and near-infrared bands quench more rapidly than SMGs. This highlights that at intermediate to low redshifts (), the majority of cosmic star formation occurs obscured by dust (Bourne et al., 2017; Dunlop et al., 2017; Bouwens et al., 2020, 2022), and particularly at , dusty star-forming galaxies contribute approximately 80% of the total cosmic star formation activity (Zavala et al., 2021).